OpenStudy (anonymous):

A $15,000 debt is to be amortized in 12 equal semiannual payments at an annual interest rate of 11% on the unpaid balance. Construct an amortization table to determine the unpaid balance after two payments have been made. Round values in the table to the nearest cent.

OpenStudy (anonymous):

does anyone know how to do these problems? I have two more.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

cannt u just use the same formula from ur last question?

OpenStudy (anonymous):

no, its different. i tried using that one.

OpenStudy (anonymous):

http://www.pine-grove.com/online-calculators/remaining-balance-calculator.htm i found this calculator but I don't know exactly what to put in

OpenStudy (abb0t):

The formula for calculating the payment amount is: \(A = P\frac{ (r(1+r)^n) }{ ((1+r)^n)-1 }\) where: A = payment Amount per period P = initial Principal (loan amount) r = interest rate per period n = total number of payments or periods

OpenStudy (anonymous):

how do I find the unpaid balance after calculating the payment amount?

OpenStudy (anonymous):

I got 2310.41, does anyone know what to do next?

OpenStudy (anonymous):

@abb0t is 2310.41 the unpaid balance?

jimthompson5910 (jim_thompson5910):

The interest rate per period is r = 0.11/2 = 0.055 There are n = 12 payments total and you're borrowing L = 15000 dollars ------------------------------------------------------- Plug all of this into the formula below P = L(r(1 + r)^12)/((1 + r)^12 - 1) P = 15000(0.055(1 + 0.055)^12)/((1 + 0.055)^12 - 1) P = 1740.43846787421 P = 1740.44 So that's the monthly payment

jimthompson5910 (jim_thompson5910):

that should give you enough info to find the remaining balance after 2 payments have been made

OpenStudy (anonymous):

11519.12?

jimthompson5910 (jim_thompson5910):

this is what I get

jimthompson5910 (jim_thompson5910):

after using this calculator http://www.pine-grove.com/online-calculators/remaining-balance-calculator.htm

OpenStudy (anonymous):

okay, I got my attempted answer by multiplying 11519 by 2 and subtracting from 15000.

jimthompson5910 (jim_thompson5910):

how did you get 11519

jimthompson5910 (jim_thompson5910):

and that seems awfully high for a payment

OpenStudy (anonymous):

i didnt pass the skills test so it doesnt tell me which answers i got wrong

jimthompson5910 (jim_thompson5910):

oh i'm sry to hear that probably got this one wrong and the one before it wrong...hmm idk